The inverse demand function for grapes is described by the equation p = 518 - 5q, where p is the price in dollars per crate and q is the number of crates of grapes demanded per week. When p = $38 per crate, what is the price elasticity of demand for grapes?
Answer #1

Answer

Elasticity of demand = (dq/dp)(p/q)

Inverse demand is given by :  

p = 518 - 5q => q = (1/5)(518 - p) => dq/dp = -1/5

It is given that p = 38 => q = q = (1/5)(518 - p) = q = (1/5)(518 - 38) = 96

Thus, we have p = 38 , q = 96 and dq/dp = -1/5

So, Elasticity of demand = (dq/dp)(p/q) = -(1/5)(38/96) = -0.0792

Hence, the price elasticity of demand for grapes = -0.0792